You want to buy a house and will need to borrow $235,000. The interest rate on your loan is 5.65 percent compounded monthly and the loan is for 20 years. What are your monthly mortgage payments?
Answer #1

This question requires application of PV of annuity formula

PV =

1- (1+ г)-т Р т Р— Рeriodic Раутent т 3 гate рer period п 3D питber of periods

n = 20 * 12 = 240 months, r = 5.65%/12 = 0.4708% (monthly)

240 (10.004708) 235, 000 P* [- 0.004708

235,000 = P * 143.5984

P = $1,636.51